OpenStudy (anonymous):

The length of a rectangle is 5/6 ft. The width is 3/8 ft. How much longer is the length of the rectangle to the width? Express your answer in lowest terms. @iGreen @PaulaLovesSchool13 @baking4lyf please help Medal!!!

OpenStudy (anonymous):

Hey Meg, what grade you in?

OpenStudy (anonymous):

7th

OpenStudy (anonymous):

Gotcha

OpenStudy (anonymous):

Pssstttt.... http://www.helpingwithmath.com/by_subject/geometry/geo_area.htm

OpenStudy (igreen):

It wants you to subtract. \(\dfrac{5}{6} - \dfrac{3}{8}\) Least Common Denominator is 24. Convert to Denominators with 24: \(\dfrac{5 \times 4}{6 \times 4} - \dfrac{3 \times 3}{8 \times 3} \rightarrow \dfrac{20}{24} - \dfrac{9}{24}\) Can you subtract 20 - 9?

OpenStudy (anonymous):

11/24?

OpenStudy (anonymous):

@iGreen am I right?

OpenStudy (igreen):

Yep, that's right.